# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

### v3

@@ Abstract
On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead; This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion). Part II introduces execution-format comparators (e.g., IEEE-754), a drift ledger, and a decision-invariance theorem that supports commensurate-accuracy claims in floating arithmetic (and that cleanly accounts for any staged/truncated auxiliary drift). Part III provides case-study reductions (LUP/solve/det/inv, LDL, blocked QR, SOI/SPD functions, GSEVP, dense LP/SDP IPM kernels, Gaussian process regression, and representative semiring problems) showing how to export the kernel advantage without reintroducing uncontrolled intermediate growth. Part IV abstracts admissible packings and extractors via a master condition and an easily checkable BWBM sufficient condition, and sketches extensions to multilinear/multigraded kernels and non-rounding extractors (e.g., CRT and semiring bucket projections).
@@ 3.1 Dyadic lattice and scaled coefficient lattice
Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ (Definition 3.1). At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that all dot-product terms $A^{\prime}_{ir}B^{\prime}_{rj}$ lie in $D^{-2}2^{-\sigma}\mathbb{Z}$.
Write each stored scalar in mantissa/exponent-offset form $x=(w/D)2^{e}$ (Appendix B.1) and denote its exponent offset by $\exp(x):=e$, with the convention $\exp(0):=+\infty$. Define for each inner index $r\in\{1,\dots,m\}$:
@@
$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}),$ (3)
and let:
$\mathcal{R}~:=~\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}.$ (4)
If $\mathcal{R}=\emptyset$ (i.e., for every $r$ either $A^{\prime}_{:,r}\equiv 0$ or $B^{\prime}_{r,:}\equiv 0$, so all dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
$g_{\min}~:=~\min_{r\in\mathcal{R}}(a_{r}+b_{r}),\\ ~~~\sigma:=\max\{0,\,-g_{\min}\}.$ (5)
@@
$\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (7)
@@
$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil$ (8)
@@ 4 Rounding primitives
Let $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ denote rounding to the nearest integer with any fixed deterministic tie-breaking convention. Fix $m\in\mathbb{Z}$ and $\delta\in(0,\frac{1}{2}]$. If $r,e\in\mathbb{R}$ satisfy:
@@ 5.1 Two-round sieve on scaled lattices
$\mathcal{S}_{\beta,\sigma}(z)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil-\beta\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z/\beta\right\rceil.$ (15)
@@ 5.3 Worked micro-example (sanity check)
so Lemma 5.2 applies and $\mathcal{S}_{\beta}(z)=T$.
@@ 6.1 Dyadic $\ell_{2}$ normalization without square roots
    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,\,K(x)-t\},$ (43)
    where $t:=\max\{0,-\min_{r}e_{r}\}$ is the shift from Definition 6.1 satisfying $S(x)=2^{-2t}\widetilde{S}(x)$. Equivalently, $k(x)=\max\{0,\,\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
@@ 7 Sufficient-base theorem
$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)\;\leq\;\frac{1}{2}-\delta_{0},$ (55)
where $\sigma$ is the node-local coefficient-lattice exponent (Definition 3.2). For dyadic bases $\beta=2^{b}$ with $b\in\mathbb{Z}_{\geq 1}$, a fully deterministic choice with no tuning is: start with the candidate $b=t+3$ where $t:=\lfloor\log_{2}(D^{2}2^{\sigma})\rfloor$; if Eq. 55 fails, increment $b$ once (doubling $\beta$). This rule (and the corresponding exact integer certificate) is spelled out in Appendix  B.5.
@@
This is Theorem  7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix  B.5. ∎
@@
Condition (54) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$. There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix  B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node. Here $\sigma$ is the node-local coefficient-lattice exponent from Definition 3.2, i.e., it is determined by the smallest dyadic lattice step on which the normalized dot-product coefficients at that node actually lie. In worst-case analyses one may upper bound $\sigma$ by a quantity proportional to recursion depth (i.e., $O(\log n)$ in a $2\times 2$ recursion), but implementations MUST compute $\sigma$ from exponent-offset metadata of $A^{\prime},B^{\prime}$ rather than setting $\sigma:=$ (recursion depth). Consequently, the required bit-width of the dyadic base $\beta=2^{b}$ grows at most logarithmically in $n$ in the worst case. This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.5). The algorithm exchanges the uncontrolled magnitude growth of classical methods for a strictly controlled, logarithmic growth in precision (staging depth), giving the soft-quadratic complexity bound.
@@ 8.1 Centering correction identity
  • Centering is separate from normalization. Centering enforces mean constraints (row and/or column means) and is inverted by the bilinear identity (Lemma 8.1). Normalization enforces $\ell_{2}$ bounds and is inverted by applying the stored dyadic exponents, i.e., are treat as separate ordered steps: center first, then normalize.
@@ 8.2 Implementation contract (integral kernel)
    Dyadic means are exact: Node dimensions are powers of two. Division by $m=2^{t}$ is implemented as an exponent-offset update.
@@
    Sufficient base with computed $\sigma$ (no depth substitution). At each node, $\sigma$ MUST be computed from the exponent offsets of the normalized operands $A^{\prime},B^{\prime}$ as in Definition 3.2 (equivalently Eq. 266). Implementations MUST NOT set $\sigma$ equal to recursion depth (this can over-inflate $\beta$ and destroy performance). Given the computed $\sigma$, implementations MUST use the fixed slack convention $\delta_{0}:=1/8$ and set the dyadic packing base $\beta=2^{b}$ deterministically by Appendix B.5. A reference implementation MUST certify the sufficient-base inequality Eq. 263 for the chosen $\beta$. If it fails at the initial candidate, apply the one-step upgrade rule $b\leftarrow b+1$ (doubling $\beta$) and re-evaluate. If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix  B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node. (Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm  2.)
@@ 8.3 Kernel pseudocode
Input: $A,B\in\mathcal{L}_{D}^{n\times n}$, with $n$ a power of two
Output: $C=AB$ exactly
@@
 return ABAB
(1) Center. Compute the exact column-mean vector $\alpha\in\mathcal{L}_{D}^{n}$ of $A$ and exact row-mean vector $\rho\in\mathcal{L}_{D}^{n}$ of $B$ using dyadic division:
$\alpha_{j}$ $:=$ $\frac{1}{n}\sum_{i=1}^{n}A_{ij},$
$\rho_{i}$ $:=$ $\frac{1}{n}\sum_{j=1}^{n}B_{ij}.$
Set $A_{0}:=A-\mathbf{1}\alpha^{\top}$ and $B_{0}:=B-\rho\mathbf{1}^{\top}$.
(2) Dyadic $\ell_{2}$ normalize (sqrt-free). For each row $i$ of $A_{0}$, compute $k_{i}:=k(A_{0,i,:})$ via Definitions 6.16.2, and scale:
$A^{\prime}_{i,:}~:=~2^{-k_{i}}A_{0,i,:}.$
For each column $j$ of $B_{0}$, compute $\ell_{j}:=k(B_{0,:,j})$ and scale:
$B^{\prime}_{:,j}~:=~2^{-\ell_{j}}B_{0,:,j}.$
@@
(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes). Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute:
$a_{r}$ $:=$ $\min_{i}\exp(A^{\prime}_{ir}),~b_{r}:=\min_{j}\exp(B^{\prime}_{rj}),$ (69)
$\mathcal{R}$ $:=$ $\{\,r:a_{r}<\infty~\text{and}~b_{r}<\infty\,\}.$ (70)
If $\mathcal{R}=\emptyset$, set $\sigma:=0$; otherwise set $g_{\min}:=\min_{r\in\mathcal{R}}(a_{r}+b_{r})$ and $\sigma:=\max\{0,-g_{\min}\}$. This is exactly Definition 3.2 / Eq. 266, and can be accumulated during Step (2).
(3) Pack, recurse, and extract quadrants. Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
/* Certified base selection (no tuning knobs; at most one upgrade). */
Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 270. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$
Evaluate the exact integer certificate Eq. 271 (and, in a reference implementation, assert the normalization bounds from Step (2))
@@
 $b\leftarrow b_{0}+1$$\beta\leftarrow 2^{b}$
 // one-step upgrade, Eq. 272
   Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation)
@@
    return ClassicalMM(A,B)(A,B)
    // exact fallback; Appendix B.5
foreach $(i,j)\in\{1,2\}^{2}$ do
   Form packed blocks (Definition 5.6):
$X_{ij}$ $:=$ $A^{\prime}_{i1}+\beta^{-1}A^{\prime}_{i2},$
$Y_{ij}$ $:=$ $B^{\prime}_{1j}+\beta B^{\prime}_{2j}.$
$Z_{ij}\leftarrow\textsc{KernelMM}(X_{ij},Y_{ij})$ $T_{ij}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{ij})$
 // Entrywise sieve (Definition 5.1)
@@
(4) Undo normalization and un-center. Undo the dyadic scalings to recover $A_{0}B_{0}$ from $T$:
$(A_{0}B_{0})_{ij}~=~2^{k_{i}+\ell_{j}}T_{ij}~~~\text{(entrywise)},$
and then recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
$C~:=~A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}.$
return $C$
Algorithm 1 KernelMM(A,B)(A,B) (dyadic lattice semantics)
@@
    $\mathcal{S}_{\beta}(Z_{ij})~=~\mathcal{S}_{\beta}(X_{ij}Y_{ij})=T_{ij},$ (73)

### v4

@@ Abstract
On fixed-width hardware, exact evaluation of the packed recursion generally requires either an extended-precision path or digit-band (slice) staging that emulates $b(n)$ bits of mantissa precision using $w$-bit words, incurring only polylogarithmic overhead. This leads to a soft-quadratic bit cost $\tilde{O}(n^{2})$ (e.g., $\Theta(n^{2}\log^{2}n)$ when $b(n)=\Theta(\log n)$ in the basic $2\times 2$ recursion). Part II introduces execution-format comparators (e.g., IEEE-754), a drift ledger, and a decision-invariance theorem that supports commensurate-accuracy claims in floating arithmetic (and that cleanly accounts for any staged/truncated auxiliary drift). Part III provides case-study reductions (LUP/solve/det/inv, LDL, blocked QR, SOI/SPD functions, GSEVP, dense LP/SDP IPM kernels, Gaussian process regression, and representative semiring problems) showing how to export the kernel advantage without reintroducing uncontrolled intermediate growth. Part IV abstracts admissible packings and extractors via a master condition and an easily checkable BWBM sufficient condition, and sketches extensions to multilinear/multigraded kernels and non-rounding extractors (e.g., CRT and semiring bucket projections).
@@ 3.1 Dyadic lattice and scaled coefficient lattice
Fix $D\in\mathbb{N}$ and recall $\mathcal{L}_{D}=D^{-1}\mathbb{Z}[1/2]$ (Definition 3.1). At each recursion node (dimension $m$), after centering and dyadic $\ell_{2}$ normalization have produced normalized operands $A^{\prime},B^{\prime}\in\mathcal{L}_{D}^{m\times m}$, we define the node-local coefficient-lattice exponent $\sigma\in\mathbb{Z}_{\geq 0}$ as the smallest integer such that every scalar product term that can appear in the middle-band $T$ and upper-band $U$ of the local three-band identity (Proposition 5.7) lies in $D^{-2}2^{-\sigma}\mathbb{Z}$. (In particular, $\sigma$ must cover the cross-half products $A^{\prime}_{i,r}B^{\prime}_{r+m/2,j}$ that contribute to the upper band $U$ under the canonical $2\times 2$ packing at this node.)
Write each stored scalar in mantissa/exponent-offset form $x=(w/D)2^{e}$ (Appendix B.1) and denote its exponent offset by $\exp(x):=e$, with the convention $\exp(0):=+\infty$. Define for each index $r\in\{1,\dots,m\}$:
@@
$b_{r}$ $:=$ $\min_{1\leq j\leq m}\exp(B^{\prime}_{rj}).$ (3)
Let:
$\mathcal{R}_{T}$ $:=$ $\{r\in\{1,\dots,m\}:a_{r}<\infty~\text{and}~b_{r}<\infty\},$
$\mathcal{R}_{U}$ $:=$ $\emptyset~\text{if}~m=1,~\text{and otherwise}~\{r\in\{1,\dots,m/2\}:a_{r}<\infty~\text{and}~b_{r+m/2}<\infty\}.$
If $\mathcal{R}_{T}=\emptyset$ and $\mathcal{R}_{U}=\emptyset$ (i.e., all relevant dot-product terms vanish), then set $\sigma:=0$. Otherwise define:
$g_{T}$ $:=$ $\min_{r\in\mathcal{R}_{T}}(a_{r}+b_{r}),$
$g_{U}$ $:=$ $\min_{r\in\mathcal{R}_{U}}(a_{r}+b_{r+m/2})~(\text{take}~g_{U}:=+\infty~\text{if}~\mathcal{R}_{U}=\emptyset),$
$g_{\min}$ $:=$ $\min\{g_{T},~g_{U}\},$
$\sigma$ $:=$ $\max\{0,~-g_{\min}\}.$
@@
$\operatorname{Round}_{\sigma}\!\left(z\right)~:=~\frac{1}{D^{2}2^{\sigma}}\left\lfloor D^{2}2^{\sigma}z\right\rceil.$ (5)
@@
$\mathcal{S}_{\beta,\sigma}(z)~:=~\operatorname{Round}_{\sigma}\!\left(z\right)-\beta\operatorname{Round}_{\sigma}\!\left(z/\beta\right)$ (6)
@@
Remark 3.3 (On “global minima” choices for $\sigma$).
A safe but overly conservative alternative is to choose $\sigma$ from the decoupled bound $\min_{i,r}\exp(A^{\prime}_{ir})+\min_{r,j}\exp(B^{\prime}_{rj})$, which allows pairing an entry of $A^{\prime}$ and an entry of $B^{\prime}$ that can never meet in any inner-product at this node. The definition above is the smallest choice that covers exactly the inner-index pairings that actually occur in the local packing (matched terms for $T$ and cross-half terms for $U$), and avoids inflating $\beta(\sigma)$ (and any staging budget) unnecessarily.
@@ 4 Rounding primitives
Let $\left\lfloor\cdot\right\rceil:\mathbb{R}\to\mathbb{Z}$ denote rounding to the nearest integer with ties-to-even. Fix $m\in\mathbb{Z}$ and $\delta\in(0,\frac{1}{2}]$. If $r,e\in\mathbb{R}$ satisfy:
@@ 5.1 Two-round sieve on scaled lattices
$\mathcal{S}_{\beta,\sigma}(z)~:=~\operatorname{Round}_{\sigma}\!\left(z\right)-\beta\operatorname{Round}_{\sigma}\!\left(z/\beta\right).$ (13)
@@ 5.3 Worked micro-example (sanity check)
so Lemma 5.2 applies and $\mathcal{S}_{\beta,0}(z)=T$.
@@ 6.1 Dyadic $\ell_{2}$ normalization without square roots
    $K(x)~=~\begin{cases}K_{0},&\text{if}~\widetilde{S}(x)\leq 2^{2K_{0}},\\ K_{0}+1,&\text{otherwise,}\end{cases}~~k(x)~=~\max\{0,~K(x)-t\},$ (41)
    where $t:=\max\{0,-\min_{r}e_{r}\}$ is the shift from Definition 6.1 satisfying $S(x)=2^{-2t}\widetilde{S}(x)$. Equivalently, $k(x)=\max\{0,~\lceil\frac{1}{2}\log_{2}\widetilde{S}(x)\rceil-t\}$, with $k(0)=0$.
@@ 7 Sufficient-base theorem
$D^{2}2^{\sigma}\left(\frac{2}{\beta}+\frac{1}{\beta^{2}}\right)~\leq~\frac{1}{2}-\delta_{0},$ (53)
where $\sigma$ is the node-local coefficient-lattice exponent (Definition 3.2). For dyadic bases $\beta=2^{b}$ with $b\in\mathbb{Z}_{\geq 1}$, a fully deterministic choice with no tuning is: start with the candidate $b=t+3$ where $t:=\lfloor\log_{2}(D^{2}2^{\sigma})\rfloor$; if Eq. 53 fails, increment $b$ once (doubling $\beta$). This rule (and the corresponding exact integer certificate) is spelled out in Appendix B.5.
@@
This is Theorem 7.1 specialized to the fixed slack $\delta=\delta_{0}$. The dyadic selection rule and the exact integer certificate are proved in Appendix B.5. ∎
@@
Condition (52) implies that $\beta$ must grow roughly in proportion to $2^{\sigma}$. There is no need for a “speed-first” policy: in exact (integral) dyadic mode we fix a small constant slack $\delta_{0}:=2^{-3}=1/8$ (not a tuning knob) and choose the dyadic base $\beta=2^{b}$ deterministically by Corollary 7.2 / Appendix B.5, with a constant-cost integer certificate that the sufficient-base inequality holds at each node. Here $\sigma$ is the node-local coefficient-lattice exponent from Definition 3.2, i.e., it is determined by the smallest dyadic lattice step on which the normalized dot-product coefficients at that node actually lie. In worst-case analyses one may upper bound $\sigma$ by a quantity proportional to recursion depth (i.e., $O(\log n)$ in a $2\times 2$ recursion), but implementations MUST compute $\sigma$ from exponent-offset metadata of $A^{\prime},B^{\prime}$ rather than setting $\sigma:=$ (recursion depth). Consequently, the required bit-width of the dyadic base $\beta=2^{b}$ grows at most logarithmically in $n$ in the worst case. This is not a violation of the model; rather, it is the precise origin of the digit-band staging factor $K(n)$ (Theorem 9.5). The algorithm exchanges the uncontrolled magnitude growth of classical methods for a strictly controlled, logarithmic growth in precision (staging depth), giving the soft-quadratic complexity bound.
@@ 8.1 Centering correction identity
  • Centering is separate from normalization. Centering enforces mean constraints (row and/or column means) and is inverted by the bilinear identity (Lemma 8.1). Normalization enforces $\ell_{2}$ bounds and is inverted by applying the stored dyadic exponents, i.e., are treated as separate ordered steps: center first, then normalize.
@@ 8.2 Implementation contract (integral kernel)
    Dyadic means are exact: All recursion node dimensions are powers of two. If the input dimension $n$ is not a power of two, pad $A,B$ with zeros to $N:=2^{\lceil\log_{2}n\rceil}$ and run KernelMM on the $N\times N$ instance (returning the top-left $n\times n$ block at the end). Division by $m=2^{t}$ is implemented as an exponent-offset update.
@@
    Sufficient base with computed $\sigma$ (no depth substitution). At each node, $\sigma$ MUST be computed from the exponent offsets of the normalized operands $A^{\prime},B^{\prime}$ as in Definition 3.2 (equivalently Eq. 289). Implementations MUST NOT set $\sigma$ equal to recursion depth (this can over-inflate $\beta$ and destroy performance). Given the computed $\sigma$, implementations MUST use the fixed slack convention $\delta_{0}:=1/8$ and set the dyadic packing base $\beta=2^{b}$ deterministically by Appendix B.5. A reference implementation MUST certify the sufficient-base inequality Eq. 286 for the chosen $\beta$. If it fails at the initial candidate, apply the one-step upgrade rule $b\leftarrow b+1$ (doubling $\beta$) and re-evaluate. If it still fails after the one-step upgrade, the implementation MUST take the failure-policy fallback of Appendix B.5 (call ClassicalMM on the original subproblem and return) and MUST NOT apply the sieve at that node. (Equivalently, in staged-precision mode an implementation may signal failure to a retry wrapper and re-attempt with larger staging; see Algorithm 2.)
@@
    Checked integer arithmetic (no silent wraparound): All mantissa arithmetic used by centering, norm accumulation, packing/unpacking, and rounding/extraction (including evaluation of the integer base certificate Eq. 296 in Appendix B.5) MUST be carried out either in a type wide enough to avoid overflow or with explicit overflow checks. Silent wraparound is not permitted. Any detected overflow (including overflow of exponent offsets) MUST be treated as leaving the declared exact-execution regime and MUST trigger the same failure policy as a failed base certificate (Appendix B.5): do not apply the sieve; instead fall back to ClassicalMM on the original subproblem and return, or signal failure to a retry wrapper (Algorithm 2).
  5. 5.
@@ 8.3 Kernel pseudocode
Algorithm 1 KernelMM(A,B)(A,B) (dyadic lattice semantics)
$A,B\in\mathcal{L}_{D}^{n\times n}$, $n\geq 1$.
$C=AB$ exactly in dyadic lattice semantics.
if $n$ is not a power of two then
  $N\leftarrow 2^{\lceil\log_{2}n\rceil}$
  Form $\widetilde{A},\widetilde{B}\in\mathcal{L}_{D}^{N\times N}$ by embedding $A,B$ in the top-left $n\times n$ block and padding all other entries with 0.
[…]
  return the top-left $n\times n$ block of $\widetilde{C}$.
end if
@@
  return $AB$.
end if
(1) Center.
For $j=1,\dots,n$, set $\alpha_{j}\leftarrow\frac{1}{n}\sum_{i=1}^{n}A_{ij}$ (exact dyadic division).
For $i=1,\dots,n$, set $\rho_{i}\leftarrow\frac{1}{n}\sum_{j=1}^{n}B_{ij}$ (exact dyadic division).
$A_{0}\leftarrow A-\mathbf{1}\alpha^{\top}$ and $B_{0}\leftarrow B-\rho\mathbf{1}^{\top}$.
(2) Dyadic $\ell_{2}$ normalize (sqrt-free).
For each row $i$ of $A_{0}$, compute $k_{i}\leftarrow k(A_{0,i,:})$ via Definitions 6.16.2, and scale $A^{\prime}_{i,:}\leftarrow 2^{-k_{i}}A_{0,i,:}$.
For each column $j$ of $B_{0}$, compute $\ell_{j}\leftarrow k(B_{0,:,j})$ and scale $B^{\prime}_{:,j}\leftarrow 2^{-\ell_{j}}B_{0,:,j}$.
@@
(2b) Compute the node-local lattice exponent $\sigma$ (no extra passes).
Using exponent offsets $\exp(\cdot)$ in the representation $x=(w/D)2^{e}$ (Appendix B.1), compute for each $r=1,\dots,n$:
  $a_{r}\leftarrow\min_{i}\exp(A^{\prime}_{ir})$ and $b_{r}\leftarrow\min_{j}\exp(B^{\prime}_{rj})$.
Define the index sets (cf. Definition 3.2):
  $\mathcal{R}_{T}\leftarrow\{r\in\{1,\dots,n\}:a_{r}<\infty~\text{and}~b_{r}<\infty\}$,
  $\mathcal{R}_{U}\leftarrow\{r\in\{1,\dots,n/2\}:a_{r}<\infty~\text{and}~b_{r+n/2}<\infty\}$.
if $\mathcal{R}_{T}=\emptyset$ and $\mathcal{R}_{U}=\emptyset$ then
  $\sigma\leftarrow 0$.
else
  $g_{T}\leftarrow\min_{r\in\mathcal{R}_{T}}(a_{r}+b_{r})$.
  $g_{U}\leftarrow\min_{r\in\mathcal{R}_{U}}(a_{r}+b_{r+n/2})$.
  $g_{\min}\leftarrow\min\{g_{T},g_{U}\}$ and $\sigma\leftarrow\max\{0,-g_{\min}\}$.
end if
This is exactly Definition 3.2 / Eq. 289 and can be accumulated during Step (2).
(3) Pack, recurse, and extract quadrants (folded across the $j$-index).
Partition $A^{\prime},B^{\prime}$ into $2\times 2$ blocks of size $(n/2)\times(n/2)$.
$\triangleright$ Certified base selection (unchanged; still certifies the inner $\beta$-sieve).
Compute the candidate exponent $b_{0}$ (i.e., $\beta_{0}=2^{b_{0}}$) from Eq. 295. Set $b\leftarrow b_{0}$ and $\beta\leftarrow 2^{b}$.
Evaluate the exact integer certificate Eq. 296 (and, in a reference implementation, assert the normalization bounds from Step (2)).
@@
  $b\leftarrow b_{0}+1$; $\beta\leftarrow 2^{b}$ $\triangleright$ one-step upgrade, Eq. 297
  Re-pack using the upgraded $\beta$ and re-evaluate the certificate (in exact integer arithmetic this second check must succeed; if it does not, treat it as overflow/truncation).
@@
   return ClassicalMM(A,B)(A,B) $\triangleright$ exact fallback; Appendix B.5
  end if
end if
$\triangleright$ Folded column packing: introduce a second dyadic base $\gamma$ to carry both block-columns at once.
Set $\gamma\leftarrow\beta^{2}$. $\triangleright$ any certified $\gamma\gg\beta$ works; $\beta^{2}$ is a simple constant-overhead choice
$\triangleright$ Single folded right embedding (both block-columns packed into $\{0,1\}$ in the $\gamma$-grade).
$Y\leftarrow\big(B^{\prime}_{11}+\beta B^{\prime}_{21}\big)~+~\gamma\big(B^{\prime}_{12}+\beta B^{\prime}_{22}\big)$.
for all $i\in\{1,2\}$ do
$\triangleright$ Left embedding (one per block-row), same as before but now independent of $j$.
  $X_{i}\leftarrow A^{\prime}_{i1}+\beta^{-1}A^{\prime}_{i2}$.
  $Z_{i}\leftarrow\textsc{KernelMM}(X_{i},Y)$. $\triangleright$ two recursive calls total
$\triangleright$ Project to the two $\gamma$-coefficients: $Z_{i}=Z_{i}^{(0)}+\gamma Z_{i}^{(1)}$ (no $\gamma^{-1}$ band).
  $Z_{i}^{(0)}\leftarrow\mathcal{S}_{\gamma,\sigma}(Z_{i})$.
  $Z_{i}^{(1)}\leftarrow(Z_{i}-Z_{i}^{(0)})/\gamma$. $\triangleright$ exact exponent shift
$\triangleright$ Inner extraction (the original $\beta$-sieve) applied to each $\gamma$-bucket.
  $T_{i1}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{i}^{(0)})$.
  $T_{i2}\leftarrow\mathcal{S}_{\beta,\sigma}(Z_{i}^{(1)})$.
end for
@@
(4) Undo normalization and un-center.
Undo the dyadic scalings to recover $A_{0}B_{0}$ from $T$ entrywise:
  $(A_{0}B_{0})_{ij}\leftarrow 2^{k_{i}+\ell_{j}}T_{ij}$.
Recover $C=AB$ using the bilinear centering identity (Lemma 8.1):
$C\leftarrow A_{0}B_{0}+\mathbf{1}(\alpha^{\top}B_{0})+(A_{0}\rho)\mathbf{1}^{\top}+(\alpha^{\top}\rho)\mathbf{1}\mathbf{1}^{\top}$.
return $C$.
@@
    $\mathcal{S}_{\beta,\sigma}(Z_{ij})~=~\mathcal{S}_{\beta,\sigma}(X_{ij}Y_{ij})=T_{ij},$ (69)